SpaceRef

SpaceRef


NASA ISS On-Orbit Status 10 July 2013 (With Video)

Status Report From: NASA HQ
Posted: Wednesday, July 10, 2013

image New Status: Post-Extravehicular Activity (EVA) Operations: Today was a half-duty work day for the USOS crew following yesterday's EVA. Flight Engineer (FE)-3 Cassidy and FE-5 Parmitano completed their post EVA health evaluation with FE-6 Nyberg acting as Crew Medical Officer (CMO). All 3 USOS crew members participated in an EVA debrief with ground teams.

USOS EVA #23 Preparations: In preparation for next Tuesday's planned USOS EVA #23, Cassidy and Parmitano completed an EVA tool configuration this morning. Additionally, based on crew request during the EVA debrief, Cassidy performed a fill of his backup Liquid Cooling and Ventilation Garment (LCVG) for use during USOS EVA #23.

Materials International Space Station Experiment (MISSE)-8 Payload Experiment Container (PEC)/ Optical Reflector Materials Experiment Ram/Wake (ORMatE-III R/W) Pack: Yesterday Parmitano taped the vent ports on the MISSE8 PEC retrieved during the EVA to minimize oxygen exposure. Today, he and Nyberg finalized packing and photography of the PEC and ORMatE for return on a future flight.

Today's Planned Tasks

Laptop RS2(1) Reboot [Complete]
SM ПСС (Caution & Warning Panel) Test [Complete]
Regeneration of БМП Ф1 Micropurification Cartridge (start) [Complete]
EKAM Battery Swap [Complete]
RSS1, RSS2 Reboot / r/g 1009 [Complete]
Daily Planning Conference (S-band) [Complete]
Removal of ССД305 Lights from ТКГ 418 (DC1) and Stowage in Spares Kit
Post-EVA Pille Dosimeter Reading / r/g 2806 [Complete]
OBSTANOVKA. Download of Obstanovka P/L Data (start) / r/g 2854 [Complete]
HAM Radio Power Up [Complete]
WRS Water Sample Analysis [Complete]
СТТС Configuration for MRM2 / Communications System [РТК] [Complete]
Periodic Health Status (PHS) Post EVA Evaluation Setup [Complete]
Set up video equipment to record ARED exercise [Complete]
URAGAN. Photo Spectrum System Battery Charge (start) / r/g 2857 [Complete]
Kulonovskiy Kristall Experiment. r/g 2861 [Complete]
System Laptop Terminal (SLT) Reboot [Complete]
Post-EVA Crew Periodic Health Status - Subject [Complete]
Post-EVA Crew Periodic Health Status (PHS) Evaluation - Crew Medical Officer (CMO) [Complete]
Photo/TV Camcorder Setup Verification [Complete]
Post-EVA Crew Periodic Health Status (PHS) Evaluation - Subject [Complete]
Post-EVA Crew Periodic Health Status (PHS) Evaluation - Crew Medical Officer (CMO) [Complete]
Periodic Health Status (PHS): Log Data and Stow Equipment [Complete]
Pumping Out Waste Water Tank [Complete]
MRM2 Comm Reconfig for Nominal Ops / Communications System [Complete]
EMU Water Refill [Complete]
KULONOVSKIY KRISTALL. Copy and Downlink Data via РСПИ / r/g 2861 [Complete]
СОЖ Maintenance r/g 2858 [Complete]
EVA Tool Config [Complete]
Photo/TV Camcorder Setup Verification [Complete]
VIZIR. Experiment Ops / r/g 2860 [Complete]
ТКГ 418 (DC1) Stowage and IMS Ops / r/g 2814 [Complete]
Reading БАОК ГАНК-4М Data [Complete]
Monitoring condition of RS structural surfaces using МВП-2К device r/g 2859 [Complete]
ИП-1 Flow Sensor Position Check r/g 2862 [Complete]
US Post-EVA Debrief [Complete]
VTR Reconfiguration [Complete]
Photo/TV Camcorder Setup Verification [Complete]
URAGAN. Observation and Photography using Photo Spectrum System / r/g 2855 [Complete]
ISERV - Opening window shutters in the LAB [Complete]
P/TV-RWS - Monitoring [Complete]
ISS - HAM Radio Session [Complete]
OBSTANOVKA. Download of Obstanovka P/L Data (termination) / r/g 2854 [Complete]
EHS Compound Specific Analyzer-Combustion (CSA-CP) Checkout [Complete]
EKAM Battery Swap [Complete]
Monitoring Hull Surfaces of RS Structural Elements using МВП-2К device. r/g 2859 [Complete]
VYNOSLIVOST. Sample removal from the panel, prep for return r/g 2863 [Complete]
MISSE8. Hardware Packing, Labeling, Photography [Complete]
TOCA 2 Data Recording [Complete]
Monitoring RS Structural Shell Surfaces Using МВП-2К Device r/g 2859 [Complete]
Photo/TV Camcorder Setup Verification [Complete]
HMS Visual Testing Activity [Deferred]
Cleaning FGB ГЖТ Detachable Screens [Complete]
P/TV-RWS - Monitoring [Complete]
IMS Update [Complete]
Photo/TV Camcorder Setup Verification [Complete]
VZAIMODEISTVIYE. Experiment Ops / r/g 2116 [Complete]
БМП Ф1 Absorption Cartridge Regeneration (termination) [Complete]
Daily Planning Conference [Complete]
EKAM Battery Swap [Complete]
REMINDER -PFE - Reading Reminder [Complete]
Task List

CEO Ops [Complete]
WHC KTO replace [Complete]
WHC EDV R&R
Roscosmos Report
Uragan Photo
BCAT C-1 Sample mix
WRS hose label
Ground

Nominal Commanding
Three-Day Look Ahead:
Thursday, 7/11: EVA DOUG review, EVA procedures review, EKAM battery swap
Friday, 7/12: ITCS samples, EVA tool config, ACE sample prep, T2 monthly inspect
Saturday, 7/13: Crew day off

QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up



// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks